OpenStudy (anonymous):

find the inequlity of 3x+4y>-12

OpenStudy (mertsj):

First subtract 3x from both sides.

OpenStudy (amistre64):

what does "find the inequality" mean? perhaps put it into y> .... form?

OpenStudy (mertsj):

Tell me what you get.

OpenStudy (anonymous):

i got y=3/4-3

OpenStudy (mertsj):

Where did the x go?

OpenStudy (amistre64):

thats good, a little bare of a few essentials; but good

OpenStudy (amistre64):

y > -3/4 x - 3

OpenStudy (anonymous):

i got the sign wrong then

OpenStudy (mertsj):

Yes. You subtracted the 3x so it should have been -3x on the right side.

OpenStudy (anonymous):

so the answer is y>-3/4x-3

OpenStudy (mertsj):

Yes. is that all you were supposed to do with it or were you supposed to graph it?

OpenStudy (anonymous):

ok i get it

OpenStudy (mertsj):

Good for you.

OpenStudy (anonymous):

\[m \le-3 or m >6\]

OpenStudy (anonymous):

hey hero help me on the this question

OpenStudy (anonymous):

\[m \le-3orm >6\]

OpenStudy (anonymous):

can you help me graph it to

hero (hero):

m \le-3\space \text{or}\space m >6 \[m \le-3\space \text{or}\space m >6\]

OpenStudy (anonymous):

yes that the question

hero (hero):

I was showing you how to write it properly using text editor

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

did know

hero (hero):